1. A wire is under a tension of 1200 N. If each meter of the wire has a mass of 10 g,
determine the velocity of the transverse wave in the wire
1
Expert's answer
2020-09-28T08:05:22-0400

The velocity of the transverse wave in the string under tension is


"v=\sqrt{\frac{T}{\mu}}=\sqrt{\frac{1200}{0.01}}=346\text{ m/s}."